### 03 Jan 2024 v1

@@ Abstract
The burgeoning field of Artificial Intelligence Generated Content (AIGC) is witnessing rapid advancements, particularly in video generation. This paper introduces AIGCBench, a pioneering comprehensive and scalable benchmark designed to evaluate a variety of video generation tasks, with a primary focus on Image-to-Video (I2V) generation. AIGCBench tackles the limitations of existing benchmarks, which suffer from a lack of diverse datasets, by including a varied and open-domain image-text dataset that evaluates different state-of-the-art algorithms under equivalent conditions. We employ a novel text combiner and GPT-4 to create rich text prompts, which are then used to generate images via advanced Text-to-Image models. To establish a unified evaluation framework for video generation tasks, our benchmark includes 11 metrics spanning four dimensions to assess algorithm performance. These dimensions are control-video alignment, motion effects, temporal consistency, and video quality. These metrics are both reference video-dependent and video-free, ensuring a comprehensive evaluation strategy. The evaluation standard proposed correlates well with human judgment, providing insights into the strengths and weaknesses of current I2V algorithms. The findings from our extensive experiments aim to stimulate further research and development in the I2V field. AIGCBench represents a significant step toward creating standardized benchmarks for the broader AIGC landscape, proposing an adaptable and equitable framework for future assessments of video generation tasks.
@@ 1 Introduction
While benchmarks for the T2V task have seen notable progress [24, 23, 19], benchmarks for the I2V task have scarcely advanced. Previous efforts like Latent Flow Diffusion Models (LFDM) [4] and CATER-GEN [17] were tested under domain-specific video scenarios. VideoCrafter [5] and I2VGen-XL [48] only utilized visual comparisons for the I2V task. Seer [11] and Stable Video Diffusion (SVD) [2] employed video-text datasets and utilized a few metrics that require reference videos. Existing I2V benchmarks suffer from 1) a lack of diverse, open-domain images with various subjects and styles to test the efficacy of different state-of-the-art algorithms; 2) an absence of a unified consensus on which evaluation metrics should be used to assess the final generated results. From the perspective of [47], these two shortcomings hinder the capability of capturing stakeholders’ concerns and interests, while also failing to construct equivalent evaluation conditions.
Figure 1: Illustration of our AIGCBench. Our AIGCBench is divided into three modules: the evaluation dataset, the evaluation metrics, and the video generation models to be assessed. Our benchmark encompasses two types of datasets: video-text and image-text datasets. To construct a more comprehensive evaluation dataset, we expand the image-text dataset by our generation pipeline. Additionally, for a thorough evaluation of video generation models, we introduce a set of evaluation metrics comprising 11 metrics across four dimensions. These metrics include both reference video-dependent and reference video-free metrics, making full use of the benchmark we propose. We also adopted human validation to confirm the rationality of the evaluation standards we proposed.
@@
To establish a comprehensive and standardized set of evaluation metrics for video generation tasks that cater to mainstream tasks such as T2V and I2V, our AIGCBench evaluates four critical dimensions: control-video alignment, motion effects, temporal consistency, and video quality, thereby capturing every aspect of video generation. This integrated framework combines metrics that are both reference video-dependent and video-free metrics, enhancing the benchmark’s rigor without exclusively relying on video-text datasets or image-text datasets alone. We strengthen this approach by incorporating image-text datasets into our evaluations, which allows us to assess content beyond the scope of existing video-text datasets and add reference video-free metrics for assessment. The experimental results demonstrate that our evaluation standard correlates well with human ratings, confirming its effectiveness. Following a thorough evaluation, we present the strengths and weaknesses of each model, alongside several insightful findings, in hopes of spurring discussions that advance the I2V field.
@@
    We evaluate I2V algorithms comprehensively using both reference video-dependent and video-free metrics across four aspects and verify the validity of our proposed evaluation standard with human judgment;
@@
Table 1: Compare the features of our AIGCBench with those of others.   and indicate whether the benchmark includes the features listed in the respective columns.
Benchmark Open-Domain Video-Text Pairs Image-Text Pairs Generated Dataset # Metrics
Prompt Complexity Image Complexity
LFDM Eval [4] 3
CATER-GEN [17] 7
Seer Eval [11] 2
VideoCrafter Eval [5] -
I2VGen-XL Eval [48] -
SVD Eval [2] 5
AnimateBench [49] 2
AIGCBench (Ours) 11
@@ 2 Background and Related Work
Current video generation primarily encompasses two major tasks: Text-to-Video (T2V) and Image-to-Video (I2V). Given the high relevance of T2V tasks to I2V tasks, we will introduce related benchmarks for T2V in Section 2.1, describe the existing evaluations for I2V tasks in Section 2.2, and briefly introduce models for I2V in Section 2.3.
@@ 2.1 Benchmarks for Text-to-Video Generation
The task of T2V differs from I2V, as videos generated from the same text can vary widely, making it less suitable for evaluation metrics that require a reference video. For T2V tasks, the results generated by different models for the same text prompt can be quite dissimilar. However, for I2V tasks, since the image imposes certain constraints, the variation in results produced by different models is generally not as pronounced. Additionally, considering that the model’s input includes image information, the complexity of the image must also be taken into account. Our AIGCBench draws on these T2V benchmarks but differs from them in several respects: 1). We need to collect or construct images for the I2V model’s input, which requires considering the comprehensiveness of both the text prompt set and the image set. 2). Although our evaluations are similar to those of the T2V task in terms of the dimensions assessed, we need to employ new evaluation standards due to the differences between T2V and I2V tasks.
@@ 2.2 Benchmarks for Image-to-Video Generation
The open-domain I2V benchmark is currently based on two main types of evaluation data: video-text and image-text datasets. Seer [11] and SVD [2] have utilized video-text datasets and employed a limited number of metrics that require reference videos for evaluation. VideoCrafter [5] and I2VGen-XL [48] have used image-text datasets and relied solely on visual comparisons. Very recently, AnimateBench [49] was released for the purpose of evaluating I2V tasks. They also generated images using text-to-image models. However, they were limited by a small number of text prompts and a limited collection of images. At the same time, there is a lack of comprehensive evaluation metrics. Both are constrained by limited evaluation datasets and an incomplete set of assessment metrics. In this paper, we expand the image-text dataset using state-of-the-art Text-to-Image models. To ensure the complexity of the generated text prompts, we generate prompts through the combinatorial traversal of four metatypes and enhance them with the capabilities of large language models. We compare our AIGCBench with other I2V benchmarks in Table 1.
2.3 Image-to-Video Generation
Thanks to the development of diffusion models [37, 38, 15, 27, 7] and multimodal techniques [30], video generation algorithms are becoming increasingly sophisticated. Early video generation was primarily based on text-to-video approaches [16, 36, 14, 45, 3, 25, 9, 21, 8, 41, 12]. However, considering that using only text can make it challenging to intuitively depict the video scenes users want to generate, image-to-video has started to gain popularity in the video generation community.
@@ 3.2.3 Generate images and filter
In order to select high-quality image-text pairs, we filtered out the top 2000 high-quality image-text pairs based on the automatic metrics from the T2I-CompBench [18]. Some examples generated by our pipeline can be seen in the lower half of Figure 2.
@@ 4 Evaluation Metrics
Our evaluation dataset includes both video-text and image-text datasets. To conduct a comprehensive evaluation, we employe two types of assessment metrics: one that requires reference videos and another that does not. In addition, we considered previous Text-to-Video benchmarks [23, 24, 19] and have integrated to propose an evaluation standard suitable for the Image-to-Video (I2V) task, covering both types of dataset. We assess the performance of different I2V models from four aspects: control-video alignment, motion effects, temporal consistency, and overall video quality. Considering that videos generated by different algorithms have varying numbers of frames, for a standardized evaluation, we adopt the approach of extracting the first 16 frames, unless otherwise specified.
@@ 4.1 Control-video alignment
Considering that current video generation tasks primarily involve two types of inputs—a starting image and a text prompt—we introduce two evaluation metrics in the first version of our benchmark: image fidelity and text-video alignment. The image fidelity metric evaluates how similar the generated video frames are to the image input into the I2V model, especially the first frame. To assess fidelity, for the first frame of the generated video, we use metrics such as Mean Squared Error (MSE) and Structural Similarity Index Measure (SSIM) [43] to calculate the degree of preservation of the first frame. For the overall video frames, we compute the image CLIP [30] similarity between the input image and each frame of the generated video. We use MSE (First), SSIM (First), and Image-GenVideo CLIP to represent these three evaluation metrics, respectively.
@@ 5.4 Findings and Discussions
Despite the notable achievements of I2V and the rapid updates of new algorithms, there is still significant room for improvement in existing solutions. After conducting a detailed survey and evaluation of the five most advanced I2V algorithms in both academia and industry, we have made the following discoveries.
@@
The input of text in I2V tasks is also crucial. Users expect to generate reasonable and aesthetically pleasing results by combining precise textual descriptions with images. Considering that most existing solutions rely on encoders from CLIP [30] or large language models [31], it is important to recognize their limitations. The CLIP [30] model is trained on image-text pairs, while large language models [31] are trained on purely textual data, making it difficult for these text encoders to capture fine-grained temporal features. We believe that there is a need to train a large-scale cross-modal model specifically for video contexts to align video and text, thereby achieving fine-grained control over video generation and enhancing user experience.
Lone video generation
@@
Currently, the speed of video generation is relatively slow. For a 3-second video, mainstream algorithms generally require about 1 minute on a V100 graphics card. Considering that video generation scenarios are based on diffusion models [37, 38, 15, 27, 7], there are currently two main routes for speeding up the process. One is to reduce the dimensionality of the video in the latent space. For example, Stable Diffusion [2] maps the video into a latent space, roughly decreasing the size of the video by about 8 times, with only a minimal loss of video quality. The other is to improve the inference speed of the diffusion model, which is also a hot research topic in the AIGC (Artificial Intelligence Generated Content) community.

### 23 Jan 2024 v3

@@ Abstract
The burgeoning field of Artificial Intelligence Generated Content (AIGC) is witnessing rapid advancements, particularly in video generation. This paper introduces AIGCBench, a pioneering comprehensive and scalable benchmark designed to evaluate a variety of video generation tasks, with a primary focus on Image-to-Video (I2V) generation. AIGCBench tackles the limitations of existing benchmarks, which suffer from a lack of diverse datasets, by including a varied and open-domain image-text dataset that evaluates different state-of-the-art algorithms under equivalent conditions. We employ a novel text combiner and GPT-4 to create rich text prompts, which are then used to generate images via advanced Text-to-Image models. To establish a unified evaluation framework for video generation tasks, our benchmark includes 11 metrics spanning four dimensions to assess algorithm performance. These dimensions are control-video alignment, motion effects, temporal consistency, and video quality. These metrics are both reference video-based and video-free, ensuring a comprehensive evaluation strategy. The evaluation standard proposed correlates well with human judgment, providing insights into the strengths and weaknesses of current I2V algorithms. The findings from our extensive experiments aim to stimulate further research and development in the I2V field. AIGCBench represents a significant step toward creating standardized benchmarks for the broader AIGC landscape, proposing an adaptable and equitable framework for future assessments of video generation tasks. We have open-sourced the dataset and evaluation code on the project website: https://www.benchcouncil.org/AIGCBench.
@@ 1 Introduction
While benchmarks for the T2V task have seen notable progress [23, 22, 18], benchmarks for the I2V task have scarcely advanced. Previous efforts like Latent Flow Diffusion Models (LFDM) [25] and CATER-GEN [16] were tested under domain-specific video scenarios. VideoCrafter [4] and I2VGen-XL [48] only utilized visual comparisons for the I2V task. Seer [10] and Stable Video Diffusion (SVD) [2] employed video-text datasets and utilized a few metrics that require reference videos. Existing I2V benchmarks suffer from 1) a lack of diverse, open-domain images [^2] Open-domain images refer to images that cover a wide variety of subjects or topics without specific restrictions on the content or category. with various subjects and styles to test the efficacy of different state-of-the-art algorithms; 2) an absence of a unified consensus on which evaluation metrics should be used to assess the final generated results. From the perspective of [47], these two shortcomings hinder the capability of capturing stakeholders’ concerns and interests, while also failing to construct equivalent evaluation conditions.
Figure 1: Illustration of our AIGCBench. Our AIGCBench is divided into three modules: the evaluation dataset, the evaluation metrics, and the video generation models to be assessed. Our benchmark encompasses two types of datasets: video-text and image-text datasets. To construct a more comprehensive evaluation dataset, we expand the image-text dataset by our generation pipeline. Additionally, for a thorough evaluation of video generation models, we introduce a set of evaluation metrics comprising 11 metrics across four dimensions. These metrics include both reference video-based and reference video-free metrics, making full use of the benchmark we propose. We also adopted human validation to confirm the rationality of the evaluation standards we proposed.
@@
To establish a comprehensive and standardized set of evaluation metrics for video generation tasks that cater to mainstream tasks such as T2V and I2V, our AIGCBench evaluates four critical dimensions: control-video alignment, motion effects, temporal consistency, and video quality, thereby capturing every aspect of video generation. This integrated framework combines metrics that are both reference video-based and video-free metrics, enhancing the benchmark’s rigor without exclusively relying on video-text datasets or image-text datasets alone. We strengthen this approach by incorporating image-text datasets into our evaluations, which allows us to assess content beyond the scope of existing video-text datasets and add reference video-free metrics for assessment. Considering the complexity and diversity of tasks, we believe that the evaluation metrics should cover at least these four aspects. For each aspect, we aim to use both reference video-based and video-free metrics. After satisfying these categorizations, the benefits of increasing the number of metrics become marginal, while it is insufficient without covering these aspects. The experimental results demonstrate that our evaluation standard correlates well with human ratings, confirming its effectiveness. Following a thorough evaluation, we present the strengths and weaknesses of each model, alongside several insightful findings, in hopes of spurring discussions that advance the I2V field.
@@
    We evaluate I2V algorithms comprehensively using both reference video-based and video-free metrics across four aspects and verify the validity of our proposed evaluation standard with human judgment;
@@
Table 1: Compare the features of our AIGCBench with those of others.   and indicate whether the benchmark includes the features listed in the respective columns. Video-based metrics, which use reference videos, contrast with video-free metrics that don’t. Considering the difficulty of the evaluation, we are not counting the sample numbers for domain-specific benchmarks [25, 16, 10].
Benchmark Open-Domain Video-Text Pairs Image-Text Pairs Generated Dataset #Samples Metric Types # Metrics
LFDM Eval [25] - Video-based 3
CATER-GEN [16] - Video-based & Video-free 7
Seer Eval [10] - Video-based 2
VideoCrafter Eval [4] - - -
I2VGen-XL Eval [48] - - -
SVD Eval [2] 900 Video-based 5
AnimateBench [49] 105 Video-free 2
AIGCBench (Ours) 3928 Video-based & Video-free 11
@@ 2 Background and Related Work
Current video generation primarily encompasses two major tasks: Text-to-Video (T2V) and Image-to-Video (I2V). Given the high relevance of T2V tasks to I2V tasks, we discuss video generation models, with a particular focus on I2V models, in Section 2.1. We will introduce related benchmarks for T2V in Section 2.2 and describe the existing benchmarks for I2V in Section 2.3.
2.1 Video Generation Models
Thanks to the development of diffusion models [36, 37, 14, 26, 6] and multimodal techniques [29], video generation algorithms are becoming increasingly sophisticated. Early video generation was primarily based on text-to-video approaches [15, 35, 13, 44, 3, 46, 24, 8, 20, 7, 40, 11]. Most of the work is based on diffusion models [37, 14, 26, 6, 12], with some being transformer-based [15, 46]. They all rely on extensive video-text or image-text datasets to train scalable models. However, considering that using only text can make it challenging to intuitively depict the video scenes users want to generate, image-to-video has started to gain popularity in the video generation community.
@@ 2.2 Benchmarks for Text-to-Video Generation
The task of T2V differs from I2V, as videos generated from the same text can vary widely, making it less suitable for evaluation metrics that require a reference video. For T2V tasks, the results generated by different models for the same text prompt can be quite dissimilar. However, for I2V tasks, since the image imposes certain constraints, the variation in results produced by different models is generally not as pronounced. This allows us to conduct a comprehensive evaluation of different Image-to-Video (I2V) algorithms on video-text datasets using evaluation metrics that are based on reference videos. Our AIGCBench draws on these T2V benchmarks but differs from them in several respects: 1). We need to collect or construct images for the I2V model’s input, which requires considering the comprehensiveness of both the text prompt set and the image set. 2). Although our evaluations are similar to those of the T2V task in terms of the dimensions assessed, we need to employ new evaluation standards due to the differences between T2V and I2V tasks.
@@ 2.3 Benchmarks for Image-to-Video Generation
The open-domain I2V benchmark is currently based on two main types of evaluation data: video-text and image-text datasets. Seer [10] and SVD [2] have utilized video-text datasets and employed a limited number of metrics that require reference videos for evaluation. VideoCrafter [4] and I2VGen-XL [48] have used image-text datasets and relied solely on visual comparisons. Very recently, AnimateBench [49] was released for the purpose of evaluating I2V tasks. They also generated images using Text-to-Image models. However, they were limited by a small number of text prompts and a limited collection of images. At the same time, there is a lack of comprehensive evaluation metrics. Both are constrained by limited evaluation datasets and an incomplete set of assessment metrics. This leads to the evaluation datasets not being representative of all stakeholders’ concerns and interests, and there is also a lack of a unified and comprehensive consensus on evaluation. In this paper, we expand the image-text dataset using state-of-the-art Text-to-Image models. To ensure the complexity of the generated text prompts, we generate prompts through the combinatorial traversal of four metatypes and enhance them with the capabilities of large language models. We compare our AIGCBench with other I2V benchmarks in Table 1.
Generating Image-Text Dataset
While most benchmarks gather datasets from the real world, CATER-GEN [16] constructs datasets using a limited set of text prompts for specific object movement scenarios. Very recently, AnimateBench [49] utilized a limited number of manually designed text prompts and also employed Text-to-Image models to generate images. However, this approach is constrained by the simplicity of the text combinations and the limited diversity of the images. Our generation pipeline uses a text combiner to randomly generate text prompts and incorporates GPT-4 [27] to enrich the content. Simultaneously, we filter the generated results to select high-quality image-text pairs.
@@ 3.2.3 Generate images and filter
In order to select high-quality image-text pairs, we filtered out the top 2003 high-quality image-text pairs based on the automatic metrics from the T2I-CompBench [17]. Some examples generated by our pipeline can be seen in the lower half of Figure 2.
@@ 4 Evaluation Metrics
Our evaluation dataset includes both video-text and image-text datasets. To conduct a comprehensive evaluation, we employ two types of assessment metrics: one that requires reference videos and another that does not. In addition, we considered previous Text-to-Video benchmarks [22, 23, 18] and have integrated to propose an evaluation standard suitable for the Image-to-Video (I2V) task, covering both types of dataset. We assess the performance of different I2V models from four aspects: control-video alignment444”control” refers to the input signals from the user, such as text, images, and other forms of control signals., motion effects, temporal consistency, and overall video quality [^5] The code is available at https://github.com/BenchCouncil/AIGCBench.. Considering that videos generated by different algorithms have varying numbers of frames, for a standardized evaluation, we adopt the approach of extracting the first 16 frames, unless otherwise specified.
@@ 4.1 Control-video alignment
The control-video alignment measures the degree of alignment between the user’s input control signals, such as text and images, and the generated video. Considering that current video generation tasks primarily involve two types of inputs—a starting image and a text prompt—we introduce two evaluation metrics in the first version of our benchmark: image fidelity and text-video alignment. The image fidelity metric evaluates how similar the generated video frames are to the image input into the I2V model, especially the first frame. To assess fidelity, for the first frame of the generated video, we use metrics such as Mean Squared Error (MSE) and Structural Similarity Index Measure (SSIM) [42] to calculate the degree of preservation of the first frame. For all frames of the video, we calculate the similarity of CLIP (Contrastive Language–Image Pre-training) [29] embeddings between the input image and each frame of the generated video. We use MSE (First), SSIM (First), and Image-GenVideo CLIP to represent these three evaluation metrics, respectively.
@@ 5.4 Findings and Discussions
Despite the notable achievements of I2V and the rapid updates of new algorithms, there is still significant room for improvement in existing solutions. Utilizing AIGCBench, we have conducted a detailed survey and evaluation of the five most advanced I2V algorithms from both academia and industry. The comprehensive analysis facilitated by AIGCBench has led us to make the following discoveries:
@@
AIGCBench’s testing methodology has revealed that current I2V tasks often fall short in allowing users to generate content with precise textual descriptions. The benchmark’s diverse datasets and nuanced evaluation metrics have shown that while solutions based on CLIP [29] and large language models [30] are a step in the right direction, they are not fully capturing the fine-grained details that yield a high degree of control over video content. AIGCBench’s detailed feedback on these algorithms has underscored the need for a model that is specifically trained for video contexts to improve text-video alignment and to provide users with that control. Our findings suggest that integrating AIGCBench’s evaluation criteria into the development process could lead to algorithms that better align with human preferences.
Long video generation
@@
Currently, the speed of video generation is relatively slow. For a 3-second video, mainstream algorithms generally require about 1 minute on a V100 graphics card. Considering that video generation scenarios are based on diffusion models [36, 37, 14, 26, 6], there are currently two main routes for speeding up the process. One is to reduce the dimensionality of the video in the latent space. For example, Stable Diffusion [2] maps the video into a latent space, roughly decreasing the size of the video by about 8 times, with only a minimal loss of video quality. The other is to improve the inference speed of the diffusion model, which is also a hot research topic in the AIGC community.
In light of these findings from AIGCBench, it becomes evident that the benchmark not only aids in the identification of current shortcomings but also offers a structure for addressing them. AIGCBench’s comprehensive framework for evaluation, encompassing a rich variety of datasets and multi-dimensional metrics, provides a roadmap for advancing the state of I2V algorithms. By exposing algorithmic weaknesses and offering a standardized platform for comparison, AIGCBench guides researchers towards developing solutions that overcome the challenges of fine-grained control, video length, and inference speed. The insights gained from AIGCBench evaluations are instrumental in pushing the boundaries of what is possible in the field of I2V generation.